OpenStudy (anonymous):

AN URN CONTAINS 3 RED MARBELS AND 7 WHITE MARBELS. A MARBLE IS DRAWN FROM URN AND A MARBLE OF THAT COLOUR IS THEN PUT IN TO THE URN. A SECOND MARBLE IS DRAWN FROM THE URN. IF BOTH MARBELS WERE OF SAME COLOUR, WHAT IS THE PROBABILITY P THAT THEY WERE BOTH WHITE?

OpenStudy (anonymous):

ANS IS 7/8

OpenStudy (alekos):

not sure how the arrived at that answer. P(white) = 7/10 on the first draw, and would also be 7/10 on the second draw which would give you a total probability of 49/100?

OpenStudy (anonymous):

EVEN I THINK THE QUESTION IS WRONG. BT IF U CONSIDER WITHOUT REPLACEMENT, THEN THW ANSWER IS CORRECT.

OpenStudy (alekos):

without replacement it should be 7/10 x 6/9 = 7/15

OpenStudy (bibby):

Stop yelling dude

OpenStudy (anonymous):

USE BAYES THEORAM. EVENT E=2WHITES + 2REDS =7/15 + 1/15 = 8/15. THEN P = 2WHITES/E=7/15/8/15=7/8

OpenStudy (alekos):

not sure about that because the final result is two white marbles. i dont believe p(red) should be considered

OpenStudy (anonymous):

ITS BAYES THEORAM, U WILL HAVE TO CONSIDER 2REDS

OpenStudy (alekos):

i dont think that bayes theorem applies

OpenStudy (anonymous):

I WILL TRY TO FIND MORE ABOUT THIS

OpenStudy (anonymous):

Welcome to OpenStudy @anshul215 ! Here's a link to help you around OpenStudy! http://openstudy.com/study#/updates/52f5a913e4b0512c88e5ef0f